# MeSH Concept Relevance and Knowledge Evolution: A Data-driven Perspective

### v2

@@ Abstract
Background: The Medical Subject Headings (MeSH), one of the main knowledge organization systems in the biomedical domain, is constantly evolving following the latest scientific discoveries in health and life sciences. Previous research focused on quantifying information in MeSH using its hierarchical structure.
Objective: In this work, we propose a data-driven approach based on information theory and network analyses to quantify the knowledge evolution in MeSH and the relevance of its individual concepts.
Methods: Our approach leverages article annotations and their citation networks to compute the level of informativeness, usefulness, disruptiveness, and influence of MeSH concepts over time. The citation network includes the instances of MeSH concepts or MeSH headings, and the concept relevance is calculated individually. Then, this computation is propagated to the hierarchy to establish the relevance of a concept.
Results: We quantitatively evaluated our approach using changes in the MeSH terminology and showed that it effectively captures the evolution of the terminology. Moreover, we validated the ability of our framework to characterize retracted articles and showed that concepts used to annotate retracted articles differ substantially from those used to annotate non-retracted ones.
Conclusion: The proposed framework provides an effective method to rank concept relevance and can be useful in maintaining evolving knowledge organization systems.
@@
journal: Biomedical Informatics
@@ 1 Introduction
Statement of significance
Problem
Quantification of information in the Medical Subject Headings (MeSH)
focuses on the structure of the hierarchy.
What is Already Known
Different approaches have been used to analyze various features, such
as the hierarchy, bibliometric indices, and the co-occurrence of descriptors.
Despite some attempts to use instances, such as the occurrence of descriptors
in PubMed articles, instance analysis still needs to be explored in
quantifying MeSH concepts over time.
What this Paper Adds
We present an approach to measure the relevance encoded by MeSH,
considering the instantiation of its concepts. The approach includes four
aspects to quantify the concept relevance: informativeness, usefulness,
disruptiveness, and influence.
@@ 2.2.4 Influence
where the parameter $\alpha$ regulates the contribution between the eigenvector and the constant term $\beta$. The term $\beta$ is used to initialize nodes with zero out-degree. The adjacency matrix of $G$ is denoted by $A_{ij}$, which captures the edges between pairs of nodes (citations in our case). The out-degree of node $j$ is represented by $k_{j}^{out}$. Similarly to disruptiveness, the centrality $x_{i}$ of an article $i$ is propagated to the concepts it is mapped to, and the overall centrality $x_{c}$ of a code is computed by summing over all the centralities of articles that concept $c$ is mapped to divided by the number of articles in the citation network, i.e., $x_{c}=\sum_{i=1}^{n}x_{i}/m$, where $x_{i}$ is the centrality score of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. We used the implementation PageRank algorithm from the NetworKit [^4] https://networkit.github.io/ library to calculate the graph centrality, where $\alpha=0.85$ and $beta=1-\alpha$.
@@ 3.1 MeSH concept relevance trends
Figure 2b shows how the first level of concepts in MeSH varies over time for each relevance aspect. All of the categories show a decreasing trend ($p-value<0.001$). The most useful concepts (Figure 2b - bottom right) in the first level are D (Chemicals and Drugs), followed by N (Health Care) and C (Diseases), indicating they encode the basic level categories of PubMed. In influence, all categories show a decreasing trend ($p-value<0.001$). Concept C is the most influential (Figure 2b - top right), indicating that ‘diseases’ is a central concept in PubMed. It is followed by the concepts D and G (Phenomena and Processes), which include biological, chemical, and physiological phenomena descriptors. There is a group of non-medical concepts at the bottom in the trend of influence, grouping categories I (Anthropology, Education, Sociology, and Social Phenomena), N, H (Disciplines and Occupations), K (Humanities), M (Named Groups), L (Information Science), J (Technology, Industry, and Agriculture), Z (Geographicals) which present a more linear trend near 0. Trends of categories in disruption are mainly decreasing ($p-value<0.001$), though categories M and H are not statistically significant. J shows an increasing trend ($p-value<0.001$). On the other hand, K presents a growing trend, and I present no trend with no statistically significant $p-value$. Concept C is also the most disruptive in the first level of the MeSH hierarchy (Figure 2b - top left), being more than two times more disruptive than the top-2 concept (D). On the other hand, non-medical categories I, N, H, K, M, L, J, and Z are located at the bottom of this figure in an almost linear trend near 0, similar to the influence aspect. Trends of entropy are primarily increasing with a $p-value<0.001$. However, categories H, G, B, A, and D present a decreasing trend ($p-value<0.001$). Category C shows no trend (not statistically significant). Concept D has the highest informativeness score (Figure 2b - bottom left) followed by concept E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment). However, it gradually decreases, with a standard deviation of 0.003377. The trends of this aspect remain constant over time, but this is due to the range of values of each MeSH category.
@@ 5 Conclusions
In this paper, we propose a methodology based on information theory and network analysis to compute the relevance of MeSH concepts for the biomedical corpus indexed by PubMed. By considering multiple relevance aspects of a concept in a data-driven analysis, our approach provides insights into the relevance evolution of MeSH concepts to the PubMed corpus. In our study, we analyzed the dynamics of MeSH over time and demonstrated the potential of our approach in various tasks, such as concept evolution and identifying retracted papers. We recognize the need for further research to expand our approach to other domains and address the computational complexities associated with some metrics. In the future, we will extend the proposed approach to different domains.

### v4

@@ Abstract
The Medical Subject Headings (MeSH), one of the main knowledge organization systems in the biomedical domain, constantly evolves following the latest scientific discoveries in health and life sciences. Previous research focused on quantifying information in MeSH using its hierarchical structure. In this work, we propose a data-driven approach based on information theory and network analyses to quantify the relevance of MeSH concepts. Our approach leverages article annotations and their citation networks to compute informativeness, usefulness, disruptiveness, and influence of MeSH concepts over time. Using the the citation network and the MeSH hierarchy, different relevance aspects are computed, and an aggregation algorithm is used to propagate the relevance scores to the parent nodes. We evaluated our approach using changes in the terminology and showed that it effectively captures the evolution of MeSH concepts. At the first level of the hierarchy, the most relevant concept - Chemical and Drugs - had a decreasing trend (p-value $<0.01$), while at the second level, the most relevant concept - Neoplasms - had an increasing trend (p-value $<0.01$). We show that the mean relevance of evolving concepts is higher for concepts that remained unchanged (2.09E-03 vs. 8.46E-04). Moreover, we validated the ability of our framework to characterize retracted articles and showed that concepts used to annotate retracted articles (mean relevance: 0.17) differ substantially from those used to annotate non-retracted ones (mean relevance: 0.15). The proposed framework provides an effective method to rank concept relevance and can be useful in maintaining evolving knowledge organization systems.
@@
journal: Journal of Knowledge-Based Systems
@@ 2.2.4 Influence
where the parameter $\alpha$ regulates the contribution between the eigenvector and the constant term $\beta$. The term $\beta$ is used to initialize nodes with zero out-degree. The adjacency matrix of $G$ is denoted by $A_{ij}$, which captures the edges between pairs of nodes (citations in our case). The out-degree of node $j$ is represented by $k_{j}^{out}$. Similarly to disruptiveness, the centrality $x_{i}$ of an article $i$ is propagated to the concepts it is mapped to, and the overall centrality $x_{c}$ of a code is computed by summing over all the centralities of articles that concept $c$ is mapped to divided by the number of articles in the citation network, i.e., $x_{c}=\sum_{i=1}^{n}x_{i}/m$, where $x_{i}$ is the centrality score of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. We used the implementation PageRank algorithm from the NetworKit [^4] https://networkit.github.io/ library to calculate the graph centrality, where $\alpha=0.85$ and $\beta=1-\alpha$.
@@ 3.1 MeSH concept relevance trends
Figure 2b shows how the first level of concepts in MeSH varies over time for each relevance aspect. All of the categories show a decreasing trend ($p-value<0.001$). The most useful concepts (Figure 2b - bottom right) in the first level are D (Chemicals and Drugs), followed by N (Health Care) and C (Diseases), indicating they encode the basic level categories of PubMed. In influence, all categories show a decreasing trend ($p-value<0.001$). Concept C is the most influential (Figure 2b - top right), indicating that ‘diseases’ is a central concept in PubMed. It is followed by the concepts D and G (Phenomena and Processes), which include biological, chemical, and physiological phenomena descriptors.
There is a group of non-medical concepts at the bottom in the trend of influence, grouping categories I (Anthropology, Education, Sociology, and Social Phenomena), N, H (Disciplines and Occupations), K (Humanities), M (Named Groups), L (Information Science), J (Technology, Industry, and Agriculture), Z (Geographicals) which present a more linear trend near 0. Trends of categories in disruption are mainly decreasing ($p-value<0.001$), though categories M and H are not statistically significant. J shows an increasing trend ($p-value<0.001$). On the other hand, K presents a growing trend, and I present no trend with no statistically significant $p-value$. Concept C is also the most disruptive in the first level of the MeSH hierarchy (Figure 2b - top left), being more than two times more disruptive than the top-2 concept (D). On the other hand, non-medical categories I, N, H, K, M, L, J, and Z are located at the bottom of this figure in an almost linear trend near 0, similar to the influence aspect. Trends of entropy are primarily increasing with a $p-value<0.001$. However, categories H, G, B, A, and D present a decreasing trend ($p-value<0.001$). Category C shows no trend (not statistically significant). Concept D has the highest informativeness score (Figure 2b - bottom left) followed by concept E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment). However, it gradually decreases, with a standard deviation of 0.003377. The trends of this aspect remain constant over time, but this is due to the range of values of each MeSH category.
@@ 5 Conclusions
In this paper, we propose a methodology based on information theory and network analysis to compute the relevance of MeSH concepts for the biomedical corpus indexed by PubMed. By considering multiple relevance aspects of a concept in a data-driven analysis, our approach provides insights into the relevance evolution of MeSH concepts to the PubMed corpus.
In our study, we analyzed the dynamics of MeSH over time and demonstrated the potential of our approach in various tasks, such as concept evolution and identifying retracted papers. We recognize the need for further research to expand our approach to other domains and address the computational complexities associated with some metrics. In the future, we will extend the proposed approach to different domains.